A point charge 4.60 μC is held fixed at the origin. A second point charge 1.20 μC with mass of 2.80×10−4 kg is placed on the x axis, 0.220 m from the origin.

What is the electric potential energy U of the pair of charges? (Take U to be zero when the charges have infinite separation.)

U

U

=
0.226

  J  

The second point charge is released from rest. What is its speed when its distance from the origin is 0.600 m .

What is its speed when its distance from the origin is 6.00 m .

What is its speed when its distance from the origin is 60.0 m .
Answer #1

Solution:

a)

U = k q1 q2 / d

U = 9 x 109 x 4.6 x 10-6 x 1.2 x 10-6 / 0.22 = 0.223 J

U = 0.226 J.

b) U = 9 x 109 x 4.6 x 10-6 x 1.2 x 10-6 / 0.6

=> U = 0.0.0828

=> 0.5 m v^2 = 0.226 - 0.0828

=> v = 31.98 m/s

c) U = 9 x 109 x 4.6 x 10-6 x 1.2x 10-6 / 6

=> U = 0.00828 J

=> 0.5 mv^2 =0.226 - 0.00828

=> v = 39.44 m/s

d) U = 9 x 109 x 4.6 x 10-6 x 1.2 x10-6 / 60

=> U = 0.000828

=>0.5 mv^2 = 0.226 - 0.000828

=> v = 40.10 m/s

thanks